OpenStudy (anonymous):

A roll of wallpaper covers an area of 28 ft2. If the roll is 1 17/24 ft wide, how long is the roll? Explain.

OpenStudy (anonymous):

28=1 17/24 * long long = 28/(1 17/28) long=16 16/41

OpenStudy (anonymous):

Is that 28 ft to the second power?

OpenStudy (anonymous):

nope ft^2 just unit, its area

OpenStudy (anonymous):

Oh, thank you! lol